# Analyze Feature Flow to Enhance Interpretation and Steering in Language Models

### v2

@@ 2.1 Linear representation hypothesis
To understand how models encode and process the information they learn, one can examine the geometric structure of their hidden representations and weights. Research has shown (Mikolov et al., 2013; Marks & Tegmark, 2024; Gurnee & Tegmark, 2024; Engels et al., 2024) that linear directions carry semantically meaningful information and may be used by models to represent learned concepts. Observations of this kind led to the development of the linear representation hypothesis, which can be stated as follows.
Hidden states $\mathbf{h}\in\mathbb{R}^{d}$ can be represented as sparse linear combinations of features $\mathbf{f}\in\mathbb{R}^{d}$ that lie in linear subspaces $\mathbb{F}\subset\mathbb{R}^{d}$. The impact of each feature is encoded by its magnitude $\|\mathbf{f}\|$. The total number of these linear subspaces with unique semantics greatly exceeds $d$, forcing the model to build an overcomplete basis in the feature space embedded in $\mathbb{R}^{d}$. During a forward pass, the model typically uses only a small fraction of them. These subspaces are usually one-dimensional lines, but more complex structures can appear (Engels et al., 2024). [^1] There is a distinction between the weak and strong LRH. The strong version posits that there are only linear representations, while the weak version says that representations are mostly linear and one-dimensional.
@@ 2.2 SAE and Transcoders
Top-K (Makhzani & Frey, 2014), allows one to control the desired level of sparsity in an SAE by fixing $k$:
@@ 3.2 Feature matching
Several methods exist for matching features between layers and modules. One approach uses correlations between activations (Wang et al., 2024; Balcells et al., 2024), but it requires considerable data to compute activation statistics. Another is a data-free approach based on SAE weights (Dunefsky et al., 2024; Balagansky et al., 2024). We found that cosine similarity between decoder weights is a valuable similarity metric, and we focus on this approach.
@@ 3.3 Tracking the evolution of feature
Currently, individual SAE features or their groups (Engels et al., 2024) are treated as units for study. However, we believe that these flow graphs may also become a compelling area for future research.
@@ 3.4 Identification of linear feature circuits
Model behavior can be decomposed into computational subnetworks, called circuits, which perform task-specific operations (Elhage et al., 2021; Marks et al., 2024). Our method helps identify potential circuits where MLP and attention modules add or remove features in a mostly linear way. High values of $s^{(M)}$ or $s^{(A)}$ are strong indicators of these circuits. We validate this in our experiments, focusing on how a feature’s meaning evolves. Examples appear in Appendix E.
@@ 4.2 Overview of experiments
Below is a concise summary of each experiment. For detailed setup, see Appendix A for the first two experiments, and Appendix B for the steering experiment.
@@
Deactivation of features.
We then measure causal relationships between a target feature and its matched predecessor by directly intervening in the hidden state. Specifically, if removing the activated predecessor from the corresponding hidden state will deactivate the target feature, we infer a causal link. We test several matching strategies and quantify effectiveness using a successful deactivation rate and an activation change score (see Appendix A).
@@
Finally, we investigate whether activating or deactivating matched features across multiple layers can enforce or prevent the generation of certain themes. We identify relevant features for a selected topic, apply multi-layer interventions to the hidden states, and measure the resulting text quality.
@@ 5.1 Identification of feature predecessors
We observe differences between datasets in the latter layers. The Python code dataset contains the least amount of natural language, and TinyStories, otherwise, has the most natural and simple language structure. The rarity of groups with activated attention could stem from our SAE training rather than an inherent property of Gemma. However, in the LLama Scope case (Figure 18), we observe a slightly similar pattern, which indicates that this is indeed the property they share.
@@ 5.2 Deactivation of features
Figure 7 further shows that deactivating a single predecessor causes a greater activation strength drop if it is a group with a single predecessor, which may indicate circuit-like behavior of combined groups.
@@ 5.3 Model steering
To test multi-layer interventions, we aim to block generation on certain topics and activate others. Our baseline is steering using the initial features from which we built flow graphs. We compare intervention on features only at particular layer $l$ (per-layer method) and from layer 00 to layer $l$ (cumulative).
We measure a Behavioral score for topic presence and Coherence for language quality. In the deactivation task, we lower the Behavioral score and define the total deactivation score as $(1-\text{Behavioral})\times\text{Coherence}$. Figure 10 shows multi-layer intervention outperforms single-layer steering of the initial feature set, reducing hyperparameter sensitivity.
Figure 9 illustrates how different $r$ values influence deactivation. Increasing $r$ shifts the best intervention layer towards the beginning of the model, while smaller $r$ allows the intervention effect to spread across many layers.
@@
Figure 10 confirms that cumulative intervention outperforms the single-layer approach, suggesting that small interventions distributed over multiple layers may be more effective.
Figure 10: Comparison of best deactivation scores. The green line indicates deactivation using only the initial feature set. Multi-layer interventions (orange, blue) perform better across different $r$ values, suggesting additional discovered features reduce hyperparameter sensitivity.
@@ 6 Discussion
We (i) confirm that $\operatorname{top}_{1}$ similarity provides valuable information about causal dependencies, and (ii) conclude that groups respond differently to the deactivation of certain predecessors, indicating that they have distinct mechanistic properties and may exhibit circuit-like behavior. The fact that residual predecessors are the most influential could be explained by the nature of the residual stream as the main communication channel.
@@
If we want to reduce a particular feature during evaluation, we typically adjust its magnitude. However, achieving a significant reduction may require large adjustment scales, which can alter the distribution of hidden states. Because we know which features contribute to the appearance of the feature we want to reduce, we can also adjust those. From this perspective, it is possible to make multiple smaller adjustments rather than one large one, thereby avoiding dramatic changes to the overall distribution of hidden states.
We conclude that multi-layer interventions indeed affect the model more than single-layer approaches in achieving the desired outcome. Flow graphs may help to understand the effect of steering and identify related features, but the downstream result depends on the properties of the specific graph. We hypothesize that removing topic-related information early allows later layers to recover general linguistic information, aligning with the ability of LLMs to “self-repair” after “damaging” the information processing flow by pruning or, in our case, intervention into the structure of hidden states (McGrath et al., 2023).
Overall, our method provides a straightforward way to identify and interpret the computational graph of the model without relying on additional data; the resulting graph can then be used for precise control over the model’s internal structure. To the best of our knowledge, we are the first to use SAE features from different layers to control LLM generation. We believe that this work opens a new perspective for zero-shot steering.
@@ Appendix C Additional results for experiments
Here we report results that did not fit in the main text.
@@ C.1 Identification of feature predecessors
The “From nowhere” group is the most present among all other groups (Figure 12). This may be the consequence of sporadic activation of some features or just a matching error. The absence of groups with an attention module is most probably the consequence of our training procedure, which clearly contrasts with the distribution for Llama Scope (Figure 18).
@@ Appendix D Experiments with Llama Scope
We have also used the Llama Scope SAE pack (He et al., 2024) to evaluate our proposed approach and have found that it is well aligned with the results we observe for Gemma Scope. However, we did not perform a steering experiment or graph building, and we consider it as one of the future study directions.
For these SAEs, the results are somewhat different, although the main picture remains the same.
@@
Second, despite the uniformity, we observe that Llama Scope groups are slightly harder to separate from each other in terms of similarity scores (Figures 18 and 18), which may also be the consequence of the different architecture of the model itself or because these SAEs initially were trained with the TopK activation function.

### v3

@@ 2.1 Linear representation hypothesis
To understand how models encode and process the information they learn, one can examine the geometric structure of their hidden representations and weights. Research has shown (Mikolov et al., 2013; Marks & Tegmark, 2024; Gurnee & Tegmark, 2024; Engels et al., 2025) that linear directions carry semantically meaningful information and may be used by models to represent learned concepts. Observations of this kind led to the development of the linear representation hypothesis, which can be stated as follows.
Hidden states $\mathbf{h}\in\mathbb{R}^{d}$ can be represented as sparse linear combinations of features $\mathbf{f}\in\mathbb{R}^{d}$ that lie in linear subspaces $\mathbb{F}\subset\mathbb{R}^{d}$. The impact of each feature is encoded by its magnitude $\|\mathbf{f}\|$. The total number of these linear subspaces with unique semantics greatly exceeds $d$, forcing the model to build an overcomplete basis in the feature space embedded in $\mathbb{R}^{d}$. During a forward pass, the model typically uses only a small fraction of them. These subspaces are usually one-dimensional lines, but more complex structures can appear (Engels et al., 2025). [^1] There is a distinction between the weak and strong LRH. The strong version posits that there are only linear representations, while the weak version says that representations are mostly linear and one-dimensional.
@@ 2.2 SAE and Transcoders
Top-K (Makhzani & Frey, 2014; Gao et al., 2025), allows one to control the desired sparsity level by fixing $k$:
@@ 3.2 Feature matching
Several methods exist for matching features between layers and modules. One approach uses correlations between activations (Wang et al., 2025; Balcells et al., 2024), but it requires considerable data to compute activation statistics. Another is a data-free approach based on SAE weights (Dunefsky et al., 2024; Balagansky et al., 2024). We found that cosine similarity between decoder weights is a valuable similarity metric, and we focus on this approach.
@@ 3.3 Tracking the evolution of feature
Currently, individual SAE features or their groups (Engels et al., 2025) are treated as units for study. However, we believe that these flow graphs may also become a compelling area for future research.
@@ 3.4 Identification of linear feature circuits
Model behavior can be decomposed into computational subnetworks, called circuits, which perform task-specific operations (Elhage et al., 2021; Marks et al., 2025). Our method helps identify potential circuits where MLP and attention modules add or remove features in a mostly linear way. High values of $s^{(M)}$ or $s^{(A)}$ are strong indicators of these circuits. We validate this in our experiments, focusing on how a feature’s meaning evolves. Examples appear in Appendix E.
@@ 4.2 Overview of experiments
Below is a concise summary of each experiment. See Appendices A and B for detailed setup.
@@
Feature Deactivation.
We measure causal relationships by intervening on hidden states: if deactivating a predecessor also deactivates target feature, we infer a causal link.
Given hidden states $\mathbf{h}$ at the predecessor’s position (previous residual, MLP, or attention output), we apply transformation $\mathbf{h}\leftarrow\mathbf{h}+a(r-1)\mathbf{v}$, where $a$ is the predecessor’s activation strength, $\mathbf{v}$ its embedding, and $r$ a rescaling coefficient ($r=0$ for deactivation). We expect this to remove the feature from the hidden state, preventing further propagation.
We evaluate four matching strategies: (1) random sampling from top-5 cosine-similarity matches, (2) permutation-based (Balagansky et al., 2024), (3) $\operatorname{top}_{1}$ cosine similarity (our method), and (4) $\operatorname{top}_{5}$ cosine similarity where all five matched features must be inactive to treat this predecessor as inactive. Effectiveness is quantified via successful deactivation rate and activation change (higher when new strength approaches 0).
@@
We test whether multi-layer feature activation/deactivation can control theme generation. For a target topic, we intervene on relevant features across layers and assess text quality.
As a baseline we use initial features from which we build flow graphs. We compare single-layer (layer $l$ only) and cumulative (layers 00 to $l$) interventions, applying the same rescaling for deactivation. For activation, we add scaled embeddings. Multi-layer strategies include linear and exponential decay of steering coefficients with respect to the layer index, and constant scale for all layers (Appendix B).
We measure Behavioral (topic presence) and Coherence (language quality) scores, and use their product as final metric (for deactivation $(1-\text{Behavioral})\times\text{Coherence}$).
@@ 5.1 Identification of feature predecessors
We observe differences between datasets in the latter layers. The Python code dataset contains the least amount of natural language, and TinyStories has the most natural and simple language structure. The rarity of groups with activated attention could stem from our SAE training rather than an inherent property of Gemma. However, in the LLama Scope case (Figure 19), we observe a slightly similar pattern, which indicates that this is indeed the property they share.
@@
We have observed that group identification performance is on par with Pearson correlation-based matching methodology. The latter reduced the ”From nowhere” group presence, but did not consistently outperform our method and performed worse on out-of-distribution Python code. See more details in Appendix C.4.
@@ 5.2 Deactivation of features
Figure 8 further shows that deactivating a single predecessor causes a greater activation strength drop if it is a group with a single predecessor, which may indicate circuit-like behavior in combined groups.
To compare our method with optimal performance, we test three approaches: (1) top-1 cosine similarity matching, (2) top-1 Pearson correlation matching, and (3) an exhaustive search for maximum achievable performance. The search procedure deactivates each active predecessor feature individually, with activation change computed only for target features identified by either cosine or Pearson matching as ”From RES”, ”From MLP”, or ”From ATT” to ensure fair comparison and computational feasibility. Testing 1,894 features across two layers (each deactivated via all three methods) yields the results in Table 1, showing comparable performance between cosine and Pearson methods, additionally validating our data-free approach.
Mean AC Success rate
Top-1 Cosine 0.75 65%
Top-1 Pearson 0.74 65%
Exhaustive Search 0.83 73%
Table 1: Comparison of deactivation methods. The exhaustive search evaluates all activated predecessor features individually and reports maximum performance. The similar results between correlation-based and our data-free method validate our approach.
@@ 5.3 Model steering
To evaluate interventions based on flow graphs, we use them to suppress or activate topics in generation. Figure 10 demonstrates that our method identifies more effective steering features across layers compared to single-feature interventions on the initial feature set. The cumulative approach additionally provides two key advantages: (1) reduced sensitivity to hyperparameter choices, and (2) improved performance with smaller hidden state perturbations.
Figure 9 analyzes the impact of rescaling coefficient $r$ on deactivation effectiveness. We observe that larger $r$ values shift the optimal intervention point toward earlier layers, while smaller $r$ values distribute the intervention effect more evenly across the network depth.
@@
Figure 10 shows that cumulative intervention outperforms the single-layer approach in a low $r$ regime, suggesting that small interventions distributed over multiple layers may be more effective for controllable generation.
Figure 10: Comparison of best deactivation scores. The green line indicates deactivation using only the initial feature set. Interventions on layers detected by our method (orange, blue) perform better across different $r$ values, suggesting additional discovered features reduce hyperparameter sensitivity.
@@ 6 Discussion
We (i) confirm that $\operatorname{top}_{1}$ similarity provides valuable information about causal dependencies, and (ii) conclude that groups respond differently to the deactivation of certain predecessors, indicating that they have distinct mechanistic properties and may exhibit circuit-like behavior. The fact that residual predecessors are the most influential could be explained by the nature of the residual stream as the main communication channel, so removing the feature at module will not prevent it from further propagation if it already exists in the residual stream.
@@
If we want to reduce a particular feature at inference, we typically adjust its magnitude. However, achieving a significant reduction may require large adjustment scales, which can alter the distribution of hidden states. Because we know which features contribute to the appearance of the feature we want to reduce, we can also adjust those. From this perspective, it is possible to make multiple smaller adjustments rather than one large one, avoiding dramatic changes to the overall distribution of hidden states.
Flow graphs may help to understand the effect of steering and identify related features, but the downstream result depends on the properties of the specific graph. Overall, we conclude that they allow to find impactful features for intervention. We hypothesize that removing topic-related information early allows later layers to recover general linguistic information, aligning with the ability of LLMs to “self-repair” after “damaging” the information processing flow by pruning or, in our case, intervention into the structure of hidden states (McGrath et al., 2023).
Overall, our method provides a straightforward way to identify and interpret the computational graph of the model without relying on additional data, achieving performance similar to Pearson correlation matching; the resulting graph can then be used for precise control over the model’s internal structure. To the best of our knowledge, we are the first to use SAE features from different layers to control LLM generation. We believe that this work opens a new perspective for zero-shot steering.
@@ C.1 Identification of feature predecessors
The “From nowhere” group is the most present among all other groups (Figure 12). This may be the consequence of sporadic activation of some features or a matching error. The absence of groups with attention module is probably the consequence of our training procedure, which clearly contrasts with the distribution for Llama Scope (Figure 19).
@@ C.3 Model steering
C.4 Comparison with Pearson Correlation Baseline
While data-driven methods provide useful insights, they face challenges with sparse SAEs and low-frequency features. Our data-free approach overcomes these limitations through adjustable top-k matching, particularly advantageous in sparse activation regimes.
We evaluated Pearson correlations on 100K non-special tokens from FineWeb’s ”default” subset for features in Gemma Scope’s even layers and all layers of Pythia-70M-Deduped and GPT-2. Using an expanded sample size of 500 (Appendix A), we identified feature groups. Figure 18 presents results for the Gemma model.
Figure 18: Feature group identification comparison (Section 5.1) between $\operatorname{top}_{1}$ cosine similarity and Pearson correlation. While correlation better captures predecessors with under-trained embeddings, it exhibits stronger dataset dependence and sparsity sensitivity.
Correlation-based matching reduced the ”From nowhere” group presence and improves predecessor identification on attention module, though potential misalignment between our attention SAEs and Gemma Scope’s residual/MLP SAEs may affect quality. Results aligned closely with Llama Scope for Pythia (showing clearer attention features), while GPT-2 displayed increasing ”From nowhere” presence in later layers.
The correlation method failed to consistently outperform $\operatorname{top}_{1}$ cosine similarity, particularly on out-of-distribution Python code. Strong agreement emerged between methods for Gemma Scope and GPT-2 residual SAEs, but weaker alignment occurred for module-based SAEs, consistent with prior feature propagation studies.
@@ Appendix D Experiments with Llama Scope
We have also used the Llama Scope SAE pack (He et al., 2024) to evaluate our proposed approach and have found that it is well aligned with the results we observe for Gemma Scope. However, we did not perform a steering experiment or graph building, and we consider it as one of the future study directions. For these SAEs, the main picture remains the same.
@@
Second, despite the uniformity, we observe that Llama Scope groups are slightly harder to separate from each other in terms of similarity scores (Figures 19 and 19), which may also be the consequence of the different architecture of the model itself or because these SAEs initially were trained with the TopK activation function.